# textwidth=0.8textheight=0.9

### v1

@@ Abstract
The field of scientific machine learning, which originally utilized multilayer perceptrons (MLPs), is increasingly adopting Kolmogorov–Arnold Networks (KANs) for data encoding. This shift is driven by the limitations of MLPs, including poor interpretability, fixed activation functions, and difficulty capturing localized or high-frequency features. KANs address these issues with enhanced interpretability and flexibility, enabling more efficient modeling of complex nonlinear interactions and effectively overcoming the constraints associated with conventional MLP architectures. This review categorizes recent progress in KAN-based models across three distinct perspectives: (i) data-driven learning, (ii) physics-informed modeling, and (iii) deep operator learning. Each perspective is examined through the lens of architectural design, training strategies, application efficacy, and comparative evaluation against MLP-based counterparts. By benchmarking KANs against MLPs, we highlight consistent improvements in accuracy, convergence, and spectral representation, clarifying KANs’ advantages in capturing complex dynamics while learning more effectively. Finally, this review identifies critical challenges and open research questions in KAN development, particularly regarding computational efficiency, theoretical guarantees, hyperparameter tuning, and algorithm complexity. We also outline future research directions aimed at improving the robustness, scalability, and physical consistency of KAN-based frameworks.
@@
Scientific Machine Learning (SciML), Kolmogorov-Arnold Network (KAN), Multilayer Perceptrons (MLP), Physics-informed Neural Network (PINN), Physics-guided Neural Network (PgNN), Data-driven Model, Deep Operator Learning, Spectral Analysis
@@ 1 Introduction
Scientific Machine Learning (SciML) is rapidly transforming the landscape of scientific computing by integrating machine learning with physicochemical first principles, mathematical models, and observational data [1, 2, 3, 4, 5, 6, 7]. It enables the development of flexible, data-efficient models that capture complex and non-linear behaviors across a wide range of natural and industrial systems, with successful applications in science and engineering; e.g., fluid mechanics [8, 9, 10, 11, 12, 13, 14, 15, 16], solid mechanics [17, 18, 19], materials science [20, 21, 22, 23, 24], climate modeling [25, 26, 27, 28, 29, 30, 31], and other disciplines [32, 33, 34, 35, 36, 37]. Traditionally, the backbone of most SciML methods has been multilayer perceptrons (MLPs) [38, 39]. Despite their success and widespread use, MLP-based approaches face persistent challenges, including poor interpretability [40], limited expressivity due to fixed activation functions [41, 42, 43], and difficulty in capturing complex, nonlinear dependencies [44, 45, 46], requiring deeper networks. These challenges have sparked growing interest in Kolmogorov–Arnold Networks (KANs), a class of neural architectures inspired by the Kolmogorov–Arnold representation theorem [47, 48, 49, 50]. Research in KANs can be broadly divided into three major modeling approaches: data-driven learning, which relies solely on observational or experimental data [51, 52]; physics-informed learning, which incorporates governing equations into the training process [53, 54, 55]; and deep operator learning, which focuses on mappings between infinite-dimensional function spaces [56, 22]. Fig. 1 presents a schematic illustration highlighting the pivotal contributions in SciML that are established using MLPs or KANs.
@@
Figure 1: A schematic illustration highlighting the pivotal contributions in SciML that are established using MLPs or KANs. For MLP-based architectures, the following are listed: deep neural network (DNN) [57], physics-guided neural network (PgNN) [58], physics-informed neural network (PiNN) [2], DeepONet [59], and Fourier neural operator [60]. For KAN-based architectures, the following are listed: Kolmogorov–Arnold representation theorem, Kolmogorov–Arnold networks, KART [47, 48, 61, 62], RBF-KAN [63], ChebKAN [64], KAN [50], Wav-KAN [65], DeepOKAN [56], PIKAN [66, 53], FBKAN [54], KAN 2.0 [67], Scaled-cPIKAN [68].
@@
The Kolmogorov–Arnold Representation Theorem (KART) [47, 48, 49] establishes that any continuous multivariate function can be represented through nested compositions of continuous univariate functions. KART demonstrates how complex multidimensional relationships can be decomposed into simpler one-dimensional transformations. While KART provides a fundamental mathematical framework for function representation, its original formulation is non-constructive, the theorem proves existence but doesn’t provide practical methods for determining the required univariate functions. This limitation makes direct application in machine learning challenging. To bridge this gap, KANs were introduced as a trainable neural architecture that mirrors the functional form of KART [50, 67]. In KANs, each univarite function in the decomposition is parametrized using basis expansions with coefficients learned from data. This setup transforms the symbolic structure into a computation graph where nonlinear operations are placed along edges, and summations are carried out at nodes [69, 63, 70]. The resulting model can be interpreted as a two-layer network aligned with the original Kolmogorov–Arnold construction [71, 72, 73]. However, this shallow structure may be insufficient for learning complex mappings in practice, especially when smooth basis functions are used. To overcome this limitation, KANs have been generalized to deeper and wider configurations, extending their expressivity while preserving the core idea of dimension-wise adaptive function learning [74, 75]. KANs have since been applied to a wide range of problems in science and engineering, including data-driven discovery of physical laws [76, 77, 78], solution of partial differential equations [79, 46, 68], and learning of operators from data [56, 22]. Beyond these domains, KANs have also been explored for tasks in image processing [80, 81, 82], audio signal classification [83], and many more applications [84, 85, 86, 87, 88, 89], demonstrating the flexibility of their functional design across diverse application areas.
@@
The integration of KANs into SciML highlights a growing alignment between model architecture and the nature of the learning task, whether data-driven, physics-informed, or operator-based. In each case, the training objectives directly influence both model design and optimization dynamics. Data-driven modeling plays a central role when physical mechanisms are incomplete, unknown, or difficult to express through explicit governing equations [51, 90, 52]. In such cases, models are trained only on observational or experimental data to identify patterns linking inputs, system states, and outputs, commonly across regression, classification, and segmentation tasks. While MLPs have traditionally served as the default architecture in these tasks, their dense, black-box nature can limit interpretability and training stability. KANs, with their structured composition of univariate functions [50, 66, 64], provide more structured approximation strategies and theoretical insights into convergence, especially under limited-data regimes [46, 91, 54, 89]. When observational data is limited or physical consistency is critical, physics-informed learning provides a complementary approach by embedding known physical laws, typically a set of coupled partial differential equations (PDEs) and closure models, into the training process. This is valuable in settings where the governing equations are known but difficult to solve due to irregular geometries, high dimensionality, or sparse and noisy measurements in inverse problems [92, 93, 94, 79]. In frameworks such as Physics-Informed Neural Networks (PINNs) [4], these constraints are enforced by incorporating PDE (or other closure model) residuals into the loss function [3, 5, 95]. However, such integration often presents trade-offs involving numerical stability, expressivity, and computational efficiency [2, 96, 97, 98, 99]. Physics-Informed KANs (PIKANs) address these challenges by replacing traditional MLP backbones with the structured KAN architecture [100, 101, 102, 85, 103, 104, 105, 106, 107, 66], allowing for more control over the solution space and potentially tighter alignment with the underlying physics. Complementing both paradigms, deep operator learning has gained attention as a general framework for approximating mappings between infinite-dimensional function spaces, rather than pointwise mappings [108, 109]. Architectures such as DeepONet [59, 110] implement this through a branch–trunk decomposition to separately encode input functions and output domains, enabling learning across varying functional inputs [111, 112, 113, 114]. While early implementations relied on densely parameterized feedforward networks, extensions like DeepOKAN adopt KAN-based operators within this framework, further improving the expressiveness and training efficiency of DeepONet-like models [56, 22, 75, 115, 116, 117]. By leveraging the KAN structure at the operator level, these models reduce sample complexity and improve approximation across broad function classes, especially in multi-fidelity and physics-informed settings [59]. Together, these approaches reflect an increasing incorporation of structure and prior knowledge into model design and training, forming a continuum that balances data availability, physical insight, and computational goals.
@@
To comprehensively evaluate the capabilities of KANs in scientific machine learning, this review is structured around three major contexts: data-driven KANs, physics-informed KANs, and deep operator KANs. For each class, we examine the architectural design principles, training strategies, and recent innovations aimed at improving performance, interpretability, and scalability. We further present representative applications and provide comparative analyses against established neural solvers such as MLPs, PINNs, and DeepONets. In the final section, we outline current challenges and propose future directions, focusing on high-dimensional scalability, mesh-independent modeling, theoretical guarantees, and advanced basis selection schemes. By organizing the review in this manner, we aim to present a unified perspective on KANs, offering both conceptual clarity and practical insight for their adoption in different learning frameworks.
@@ 2 Kolmogorov-Arnold Representation Theorem
Kolmogorov’s superposition theorem, originally introduced in 1957 by Kolmogorov [47, 118] and later refined by Arnold in a series of papers [48, 49], is a fundamental result in approximation theory that addresses whether arbitrary multivariate continuous functions can be decomposed into simpler univariate functions. Kolmogorov demonstrated that any continuous real-valued function defined on an $n$-dimensional domain can be expressed as a finite composition of continuous univariate functions and the addition operation. Arnold further refined this result, particularly for the case of three variables, shedding light on the structural essence of these representations. Over the years, this theorem has been improved by several researchers, including Lorentz [119, 120], Sprecher [121, 122], and Friedman [123], among others. Hecht-Nielsen’s interpretation [124, 125, 121] showed limitations due to the non-smoothness of the inner functions and the dependency of the outer functions on the specific target function. Kůrková [126, 127] addressed some of these issues by introducing approximations using sigmoidal functions, enabling practical implementations. Sprecher [128, 129] developed a numerical algorithm for a constructive proof of the theorem, though challenges regarding the continuity and monotonicity of the inner functions persisted. Köppen [62] proposed modifications to overcome these shortcomings, and Braun and Griebel [130] advanced this work by rigorously proving the existence, continuity, and monotonicity of the modified inner functions, providing a complete constructive proof based on Sprecher’s framework with Köppen’s improvements. This theorem, often referred to as the Kolmogorov-Arnold Representation Theorem (KART), remains a cornerstone in the study of function approximation and has found applications in neural networks, despite its inherent complexities. Let $f:[0,1]^{n}\to\mathbb{R}$ be a continuous function. Then KART states that there exist continuous univariate functions $\varphi_{k}:\mathbb{R}\to\mathbb{R}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \psi_{k,j}:\mathbb{R}\to\mathbb{R},$ for indices $k=0,1,\ldots,2n$, and $j=1,2,\ldots,n$, such that,
@@
Intuitively, this means that one can build a continuous function of $n$ variables by summing together only univariate inner functions $\psi_{k,j}$ (one for each dimension) inside univariate outer functions $\varphi_{k}$. Thus, any continuous map $f$ from the $n$-dimensional space to the real line can be realized by a relatively simple set of one-dimensional building blocks and repeated additions. This canonical structure corresponds to what is referred to in KANs as the original depth-2 representation, where the first layer consists of learnable univariate functions $\psi_{k,j}$ applied independently to each input variable, and the second layer consists of univariate functions $\varphi_{k}$ composed with the linear combinations of the outputs from the first layer. In other words, a single-layer KAN, as originally inspired by KART, implements a two-stage functional composition: first applying input-wise transformations, and then aggregating them through summation and another univariate nonlinearity. This architecture forms the mathematical and conceptual foundation of KANs. The theorem is often considered a cornerstone in the theory of universal function approximation [131, 132, 133, 134], because it ensures that high-dimensional continuous functions can be decomposed into low-dimensional components. This insight underpins much of modern approximation theory [121, 130, 135], reinforcing the idea that high-dimensional phenomena may be tractably approximated using compositions of simpler (often one-dimensional) functions.
@@
Although the theorem predates modern machine learning by several decades, it provides a theoretical basis for understanding how complex functions (such as classification or regression functions in deep learning) can, in principle, be approximated by systematic compositions of simple functions. In particular, the idea that combining “depth” (stacking functions or sequential composition) and “width” (adding functions in parallel) can help approximate complex high-dimensional mappings is similar to several theoretical arguments about the universality of neural networks [136, 137, 138]. In this sense, Kolmogorov-Arnold representations show why neural architectures built from repeated layers of simple univariate transformations [125, 139, 140] may achieve remarkable expressiveness in practice. Although this result is fundamentally a theorem in classical analysis, its implications continue to inspire research in high-dimensional approximation, function representations, and computational architectures for learning complex relationships in data. Further refinements and variations of Kolmogorov-Arnold’s original ideas continue to appear in the mathematics of approximation theory [127, 130, 141], functional analysis [142, 143], and theoretical computer science [140, 144, 145].

### v2

@@ Abstract
The field of scientific machine learning, which originally utilized multilayer perceptrons (MLPs), is increasingly adopting Kolmogorov–Arnold Networks (KANs) for data encoding. This shift is driven by the limitations of MLPs, including poor interpretability, fixed activation functions, and difficulty capturing localized or high-frequency features. KANs address these issues with enhanced interpretability and flexibility, enabling more efficient modeling of complex nonlinear interactions and effectively overcoming the constraints associated with conventional MLP architectures. This review categorizes recent progress in KAN-based models across three distinct perspectives: (i) data-driven learning, (ii) physics-informed modeling, and (iii) deep-operator learning. Each perspective is examined through the lens of architectural design, training strategies, application efficacy, and comparative evaluation against MLP-based counterparts. By benchmarking KANs against MLPs, we highlight consistent improvements in accuracy, convergence, and spectral representation, clarifying KANs’ advantages in capturing complex dynamics while learning more effectively. In addition to reviewing recent literature, this work also presents several comparative evaluations that clarify central characteristics of KAN modeling and hint at their potential implications for real-world applications. Finally, this review identifies critical challenges and open research questions in KAN development, particularly regarding computational efficiency, theoretical guarantees, hyperparameter tuning, and algorithm complexity. We also outline future research directions aimed at improving the robustness, scalability, and physical consistency of KAN-based frameworks.
@@
Scientific Machine Learning (SciML), Kolmogorov-Arnold Network (KAN), Physics-informed Neural Network (PINN), Data-driven Model, Deep-operator Learning, Spectral Analysis
@@ 1 Introduction
Scientific Machine Learning (SciML) is rapidly transforming the landscape of scientific computing by integrating machine learning with physicochemical first principles, mathematical models, and observational data thiyagalingam2022scientific , raissi2019physics , karniadakis2021physics , raissi2017physicsinformeddeeplearning , cai2021physics , doi:10.1137/24M1643232 , doi:10.1137/24M1646455 . It enables the development of flexible, data-efficient models that capture complex and non-linear behaviors across a wide range of natural and industrial systems, with successful applications in science and engineering; e.g., fluid mechanics jin2021nsfnets , MAO2020112789 , JAGTAP2022111402 , faroughi2022meta , pawar2024geo , LOU2021110676 , SHU2023111972 , OVADIA2025117982 , faroughi2024physics , solid mechanics jin2023recent , haghighat2020deep , zhu2025extended , zeng2024adaptive , kumar2022machine , materials science CIESIELSKI2025113417 , datta2023multihead , kiyani2025predicting , NIELSEN2002177 , kiyani2025crack , climate modeling PARTEE2022101707 , Eyring2024 , doi:10.1098/rsta.2020.0093 , soltanmohammadi2023comparative , pawar2024esm , NEURIPS2023_45fbcc01 , ehteram2024read , mostajeran2025context , and other disciplines li2022scientific , bedolla2020machine , tapeh2023artificial , celaya2025adaptive , celaya2025learning , celaya2024solutions , yang2025pinn , maceachern2021machine , mostajeran2025solving , he2024phase , darcet2023vision , grandits2024neural . Traditionally, the backbone of most SciML methods has been multilayer perceptrons (MLPs) JAGTAP2020109136 , li2024physics . Despite their success and widespread use, MLP-based approaches face persistent challenges, including poor interpretability cranmer2023interpretable , limited expressivity due to fixed activation functions glorot2010understanding , goodfellow2016deep , raghu2017expressive , and difficulty in capturing complex, nonlinear dependencies hirsch2024neural , cruz2025state , mostajeran2024epi , requiring deeper networks. These challenges have sparked growing interest in Kolmogorov–Arnold Networks (KANs), a class of neural architectures inspired by the Kolmogorov–Arnold representation theorem kolmogorov1957representations , Arnold1958 , arnold1959representation , liu2024kan . Research in KANs can be broadly divided into three major modeling approaches: data-driven learning, which relies solely on observational or experimental data kou2021data , bahramian2023data ; physics-informed learning, which incorporates governing equations into the training process wang2024kolmogorov , howard2024finite , toscano2024inferring ; and deep-operator learning, which focuses on mappings between infinite-dimensional function spaces abueidda2025deepokan , kiyani2025predicting . Fig. 1 presents a schematic illustration highlighting the pivotal contributions in SciML that are established using MLPs or KANs.
@@
Figure 1: A schematic illustration highlighting the pivotal contributions in SciML that are established using MLPs or KANs. For MLP-based architectures, the following are listed: deep neural network (DNN) ivakhnenko1967cybernetics , physics-guided neural network (PgNN) andersen2002artificial , physics-informed neural network (PiNN) raissi2019physics , DeepONet lu2021learning , and Fourier neural operator Li2020FNO . For KAN-based architectures, the following are listed: Kolmogorov–Arnold representation theorem, Kolmogorov–Arnold networks, KART kolmogorov1957representations , Arnold1958 , sprecher2002space , koppen2002training , RBF-KAN li2024kolmogorov , ChebKAN ss2024chebyshev , KAN liu2024kan , Wav-KAN bozorgasl2405wav , DeepOKAN abueidda2025deepokan , PIKAN shukla2024comprehensive , wang2024kolmogorov , FBKAN howard2024finite , KAN 2.0 liu2024kan2 , Scaled-cPIKAN mostajeran2025scaled , KKAN toscano2025kkans .
@@
The Kolmogorov–Arnold Representation Theorem (KART) kolmogorov1957representations , Arnold1958 , arnold1959representation establishes that any continuous multivariate function can be represented through nested compositions of continuous univariate functions. KART demonstrates how complex multidimensional relationships can be decomposed into simpler one-dimensional transformations. While KART provides a fundamental mathematical framework for function representation, its original formulation is non-constructive, the theorem proves existence but doesn’t provide practical methods for determining the required univariate functions. This limitation makes direct application in machine learning challenging. To bridge this gap, KANs were introduced as a trainable neural architecture that mirrors the functional form of KART liu2024kan , liu2024kan2 . In KANs, each univarite function in the decomposition is parametrized using basis expansions with coefficients learned from data. This setup transforms the symbolic structure into a computation graph where nonlinear operations are placed along edges, and summations are carried out at nodes liu2024ikan , li2024kolmogorov , yang2024kolmogorov . The resulting model can be interpreted as a two-layer network aligned with the original Kolmogorov–Arnold construction girosi1989representation , ismailov2014approximation , schmidt2021kolmogorov , poluektov2023construction . However, this shallow structure may be insufficient for learning complex mappings in practice, especially when smooth basis functions are used. To overcome this limitation, KANs have been generalized to deeper and wider configurations, extending their expressivity while preserving the core idea of dimension-wise adaptive function learning basina2024kat , pant2025mlps . KANs have since been applied to a wide range of problems in science and engineering, including data-driven discovery of physical laws xu2024fourierkan , peng2024predictive , genet2024temporal , solution of partial differential equations shukla2020physics , mostajeran2024epi , mostajeran2025scaled , and learning of operators from data abueidda2025deepokan , kiyani2025predicting . Beyond these domains, KANs have also been explored for tasks in image processing azam2024suitability , cheon2024kolmogorov , seydi2024exploring , audio signal classification yu2024kan , and many more applications li2025u , shuai2025physics , karakacs2024novel , seydi2024unveiling , cheon2024demonstrating , panahi2025data , demonstrating the flexibility of their functional design across diverse application areas.
@@
The integration of KANs into SciML highlights a growing alignment between model architecture and the nature of the learning task, whether data-driven, physics-informed, or operator-based. In each case, the training objectives directly influence both model design and optimization dynamics. Data-driven modeling plays a central role when physical mechanisms are incomplete, unknown, or difficult to express through explicit governing equations kou2021data , williams2021evolution , bahramian2023data , ghorbani2025using , zhou2025askan , mollaali2025conformalized . In such cases, models are trained only on observational or experimental data to identify patterns linking inputs, system states, and outputs, commonly across regression, classification, and segmentation tasks. While MLPs have traditionally served as the default architecture in these tasks, their dense, black-box nature can limit interpretability and training stability. KANs, with their structured composition of univariate functions liu2024kan , shukla2024comprehensive , ss2024chebyshev , provide more structured approximation strategies and theoretical insights into convergence, especially under limited-data regimes mostajeran2024epi , ejaz2024can , howard2024finite , panahi2025data . When observational data is limited or physical consistency is critical, physics-informed learning provides a complementary approach by embedding known physical laws, typically a set of coupled partial differential equations (PDEs) and closure models, into the training process. This is valuable in settings where the governing equations are known but difficult to solve due to irregular geometries, high dimensionality, or sparse and noisy measurements in inverse problems pang2019fpinns , YANG2021109913 , shukla2021parallel , shukla2020physics . In frameworks such as Physics-informed Neural Networks (PINNs) raissi2017physicsinformeddeeplearning , these constraints are enforced by incorporating PDE (or other closure model) residuals into the loss function karniadakis2021physics , cai2021physics , kharazmi2021hp . However, such integration often presents trade-offs involving numerical stability, expressivity, and computational efficiency raissi2019physics , nabian2020adaptive , cuomo2022scientific , psaros2023uncertainty , kashefi2025physics . Physics-informed KANs (PIKANs) address these challenges by replacing traditional MLP backbones with the structured KAN architecture toscano2025pinns , jacob2024spikans , daryakenari2025representation , shuai2025physics , patra2024physics , guo2024physics , faroughi2025neural , xiapikans , kiyani2025optimizer , shukla2024comprehensive , allowing for more control over the solution space and potentially tighter alignment with the underlying physics. Complementing both paradigms, deep-operator learning has gained attention as a general framework for approximating mappings between infinite-dimensional function spaces, rather than pointwise mappings wen2022u , azizzadenesheli2024neural . Architectures such as DeepONet lu2021learning , lu2021deepxde implement this through a branch–trunk decomposition to separately encode input functions and output domains, enabling learning across varying functional inputs lu2022multifidelity , goswami2022physics , kobayashi2024improved , garg2023vb . While early implementations relied on densely parameterized feedforward networks, extensions like DeepOKAN adopt KAN-based operators within this framework, further improving the expressiveness and training efficiency of DeepONet-like models abueidda2025deepokan , kiyani2025predicting , pant2025mlps , wang2025efkan , yu2025deepoheat , toscano2025kkans . By leveraging the KAN structure at the operator level, these models reduce sample complexity and improve approximation across broad function classes, especially in multi-fidelity and physics-informed settings lu2021learning . Together, these approaches reflect an increasing incorporation of structure and prior knowledge into model design and training, forming a continuum that balances data availability, physical insight, and computational goals.
@@
Several surveys have examined the development and applications of KANs from different perspectives somvanshi2025survey , ji2024comprehensive , essahraui2025kolmogorov . For instance, ji2024comprehensive presents an analytical overview that focuses on empirical comparisons between KANs and conventional MLPs, offering a detailed discussion of architectural variants, interpretability approaches, and performance trade-offs. In addition, essahraui2025kolmogorov provides a structured overview of existing KAN architectures and their applications, emphasizing interpretability as a central feature and addressing key computational and implementation challenges. In contrast, to comprehensively evaluate the capabilities of KANs in scientific machine learning, this review is structured around three major contexts: data-driven KANs, physics-informed KANs, and deep-operator KANs. For each class, we examine the architectural design principles, training strategies, and recent innovations aimed at improving performance, interpretability, and scalability. We further present representative applications and provide comparative analyses against established neural solvers such as MLPs, PINNs, and DeepONets. Alongside reviewing prior studies, the paper also presents several comparative analyses carried out to clarify central characteristics of KAN schemes and how they might be applied in practice. In the final section, we outline current challenges and propose future directions, focusing on high-dimensional scalability, mesh-independent modeling, theoretical guarantees, and advanced basis selection schemes. By organizing the review in this manner, we aim to present a unified perspective on KANs, offering both conceptual clarity and practical insight for their adoption in different learning frameworks.
@@ 2 Kolmogorov-Arnold Representation Theorem
Kolmogorov’s superposition theorem, originally introduced in 1957 by Kolmogorov kolmogorov1957representations , kolmogorov1961representation and later refined by Arnold in a series of papers Arnold1958 , arnold1959representation , is a fundamental result in approximation theory that addresses whether arbitrary multivariate continuous functions can be decomposed into simpler univariate functions. Kolmogorov demonstrated that any continuous real-valued function defined on an $n$-dimensional domain can be expressed as a finite composition of continuous univariate functions and the addition operation. Arnold further refined this result, particularly for the case of three variables, shedding light on the structural essence of these representations. Over the years, this theorem has been improved by several researchers, including Lorentz lorentz1966approximation , lorentz1996constructive , Sprecher sprecher1965structure , sprecher1972improvement , and Friedman fridman1967improvement , among others. Hecht-Nielsen’s interpretation hecht1987counterpropagation , hecht1987kolmogorov , sprecher1965structure showed limitations due to the non-smoothness of the inner functions and the dependency of the outer functions on the specific target function. Kůrková kuurkova1991kolmogorov , kuurkova1992kolmogorov addressed some of these issues by introducing approximations using sigmoidal functions, enabling practical implementations. Sprecher sprecher1996numerical , sprecher1997numerical developed a numerical algorithm for a constructive proof of the theorem, though challenges regarding the continuity and monotonicity of the inner functions persisted. Köppen koppen2002training proposed modifications to overcome these shortcomings, and Braun and Griebel braun2009constructive advanced this work by rigorously proving the existence, continuity, and monotonicity of the modified inner functions, providing a complete constructive proof based on Sprecher’s framework with Köppen’s improvements. This theorem, often referred to as the Kolmogorov-Arnold Representation Theorem (KART), remains a cornerstone in the study of function approximation and has found applications in neural networks, despite its inherent complexities. Let $f:[0,1]^{n}\to\mathbb{R}$ be a continuous function. Then KART states that there exist continuous univariate functions $\varphi_{k}:\mathbb{R}\to\mathbb{R}\penalty 10000\ \text{and}\penalty 10000\ \psi_{k,j}:\mathbb{R}\to\mathbb{R},$ for indices $k=0,1,\ldots,2n$, and $j=1,2,\ldots,n$, such that,
@@
Intuitively, this means that one can build a continuous function of $n$ variables by summing together only univariate inner functions $\psi_{k,j}$ (one for each dimension) inside univariate outer functions $\varphi_{k}$. Thus, any continuous map $f$ from the $n$-dimensional space to the real line can be realized by a relatively simple set of one-dimensional building blocks and repeated additions. This canonical structure corresponds to what is referred to in KANs as the original depth-2 representation, where the first layer consists of learnable univariate functions $\psi_{k,j}$ applied independently to each input variable, and the second layer consists of univariate functions $\varphi_{k}$ composed with the linear combinations of the outputs from the first layer. In other words, a single-layer KAN, as originally inspired by KART, implements a two-stage functional composition: first applying input-wise transformations, and then aggregating them through summation and another univariate nonlinearity. This architecture forms the mathematical and conceptual foundation of KANs. The theorem is often considered a cornerstone in the theory of universal function approximation cybenko1989approximation , hornik1989multilayer , hornik1991approximation , barron1993universal , because it ensures that high-dimensional continuous functions can be decomposed into low-dimensional components. Moreover, ismayilova2024kolmogorov , ismailov2023three show that in the Kolmogorov two–hidden–layer neural network, the outer activation can inherit the continuity, discontinuity, and boundedness of the target function, allowing it to represent all types of multivariate functions, whether continuous, discontinuous, bounded, or unbounded. This insight underpins much of modern approximation theory sprecher1965structure , braun2009constructive , yarotsky2017error , reinforcing the idea that high-dimensional phenomena may be tractably approximated using compositions of simpler (often one-dimensional) functions.
@@
Although the theorem predates modern machine learning by several decades, it provides a theoretical basis for understanding how complex functions (such as classification or regression functions in deep learning) can, in principle, be approximated by systematic compositions of simple functions. In particular, the idea that combining “depth” (stacking functions or sequential composition) and “width” (adding functions in parallel) can help approximate complex high-dimensional mappings is similar to several theoretical arguments about the universality of neural networks telgarsky2016benefits , lu2017expressive , schmidt2020nonparametric . In this sense, Kolmogorov-Arnold representations show why neural architectures built from repeated layers of simple univariate transformations hecht1987kolmogorov , alonso2024mathematics , kilani2024kolmogorov may achieve remarkable expressiveness in practice. Although this result is fundamentally a theorem in classical analysis, its implications continue to inspire research in high-dimensional approximation, function representations, and computational architectures for learning complex relationships in data. Further refinements and variations of Kolmogorov-Arnold’s original ideas continue to appear in the mathematics of approximation theory kuurkova1992kolmogorov , braun2009constructive , igelnik2003kolmogorov , functional analysis liu2015kolmogorov , tikhomirov1963kolmogorov , and theoretical computer science kilani2024kolmogorov , borenstein2006kolmogorov , edmunds2006approximation .